write a quadratic equation that describes the area of a rectangle whose perimeter is 24 feet.
Answer by ankor@dixie-net.com(22740)

write a quadratic equation that describes the area of a rectangle whose perimeter is 24 feet.
:
2L + 2W = 24
Simplify
L + W = 12
L = (12 - W)
:
Area = L*W
Substitute (12-W) for L
A = (12-W)*W
:
A = -W^2 + 12W: a quadratic equation describing the area